### v1

# First Digit Distributions of Gamma-Ray Bursts

Abstract
The occurrence of the first significant digits from real world sources is usually not equally distributed, but is consistent with a logarithmic distribution instead, known as Benford’s law. In this work, we perform a comprehensive investigation on the first digit distributions of the duration, fluence, and energy flux of gamma-ray bursts (GRBs) for the first time. For a complete GRB sample, we find that the first digits of the duration and fluence adhere to Benford’s law. However, the energy flux shows a significant departure from this law, which may be due to the fact that a considerable part of the energy flux measurements are restricted by lack of spectral information. Based on the conventional duration classification scheme, we also check if the durations and fluences of long and short GRBs (with duration s and s, respectively) obey Benford’s law. We find that the fluences of both long and short GRBs still agree with the Benford distribution, but their durations do not follow Benford’s law. Our results hint that the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme.
keywords:
gamma-ray burst: general — methods: statistical — astronomical data bases: miscellaneous1 Introduction
People might think that the first significant digits (i.e., 1,2,…,9) of any randomly chosen dataset tend to be uniformly distributed, but it is not true in the natural world. As early as in 1881, Simon Newcomb had observed an unanticipated pattern in the first digits of logarithm tables: the number 1 appears more frequently than 2, 2 than 3, and so on (Newcomb 1881). More than a half century later, Frank Benford rediscovered that there is a logarithmic distribution of first digits in numerous data tables, which is often called the first digit law or Benford’s law (Benford 1938). This law states that for a given real database, the probability of numbers with the first digit is expressed as (Benford 1938)
Empirically, Benford’s law has been proven in various research fields, including geography (e.g., the lengths of rivers and the areas of lakes; Benford 1938), finance (e.g., the stock market indices; Ley 1996; De Ceuster et al. 1998), biology (e.g., pre-vaccination measles incidence data, absolute values from human magnetoencephalography recordings, and gene data lengths of bacteria; Cáceres et al. 2008), seismology (e.g., the recurrence times of seismic events; Sottili et al. 2012), statistical and nuclear physics (e.g., physical constants and distributions (Burke & Kincanon 1991; Shao & Ma 2010b), half-lives of unstable nuclei (Buck et al. 1993; Ni & Ren 2008; Ni et al. 2009), widths of hadrons (Shao & Ma 2009), and the lepton branching fractions (Dantuluri & Desai 2018)), etc. In practice, this peculiar law has been effectively used to distinguish and diagnose frauds in taxing and accounting (Nigrini 1996; Nigrini & Mittermaier 1997; Geyer & Williamson 2004), and to minimize storage space and speed up calculation in computer science (Barlow & Bareiss 1985; Schatte 1988; Berger & Hill 2007). Theoretically, this law has been well explained by using a central-limit-like theorem for first digits (Hill 1995) and a simple Markov process (Burgos & Santos 2021). In mathematics, Benford’s law is scale invariant (Berger et al. 2008), which indicates that it is independent of any particular choice of units (Pinkham 1961).
In astronomy, Benford’s law also has extensive application in all sorts of astrophysical datasets, such as light curves of variable stars and other X-ray sources (Moret et al. 2006), pulsar properties (Shao & Ma 2010a), distances of galaxies and stars (Alexopoulos & Leontsinis 2014), exoplanetary and asteroid data (Shukla et al. 2017; Melita & Miraglia 2021), GAIA DR2 parallaxes (de Jong et al. 2020), and so on. Nevertheless, there are also some types of data, e.g., pulsar and fast radio burst dispersion measures, which do not obey Benford’s law (Mamidipaka & Desai 2023).
In this work, we investigate the first digit distributions of the duration, fluence, and energy flux of gamma-ray bursts (GRBs), and check if these first digits conform to Benford’s law for the first time. GRBs are flashes of high-energy radiation originating from energetic explosions in the Universe. According to their duration time (the time interval observed to contain 90% of the prompt emission), GRBs can be classified into long GRBs ( s) and short GRBs ( s) (Kouveliotou et al. 1993). Generally, long GRBs are suggested to be powered by the core collapses of massive stars (Paczyński 1998; Woosley & Bloom 2006), and short GRBs by the mergers of binary compact objects (Eichler et al. 1989; Narayan et al. 1992).
2 Observational data and statistical results
2.1 Dataset
We download the durations (in units of s), fluences (in units of erg ), and energy fluxes (in units of erg ) in the 10–1000 keV energy range from the online catalog of GRBs observed by Fermi’s Gamma-ray Burst Monitor (Fermi-GBM) (Gruber et al. 2014; von Kienlin et al. 2014; Narayana Bhat et al. 2016; von Kienlin et al. 2020).111https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html The Fermi-GBM burst catalog comprises of a list of 3665 cosmic GRBs between 2008 July 12 and 2023 December 6. We remove 1 GRB for which no relative data were available. We carry out the first digit analyses for all the remaining 3664 GRBs with and fluence measurements. The energy flux of each burst in the observer frame is calculated as
where is the peak flux on the 64 ms timescale (in units of photon ), the spectral model is the Band function (Band et al. 1993), and are 10 keV and 1000 keV, respectively. It is obvious that the spectral parameters are required to calculate the energy flux. But not every burst has the spectral information. There are only 2298 GRBs with energy flux measurements in the catalog.
2.2 Results
The first digit distributions of the duration and fluence for the complete GRB sample are presented in Figure 1 and Table 1. As described above, there are totally available GRBs. The expected number according to Benford’s law, , along with the root mean square error estimated by the binomial distribution, , are also shown in the figure. From Figure 1, we can see that the observed distributions are well consistent with the theoretical predictions from Benford’s law. In order to quantify the goodness of fit of Benford’s law, we adopt the Pearson
where and are the observed number and the expected Benford number for a single digit , respectively. For the first digit distribution of the duration, we obtain a Pearson value of 13.0 for 8 degrees of freedom. For the fluence distribution, we obtain a value of 12.5. These two values correspond to -values of 0.11 and 0.13, respectively, which strongly support the null hypothesis that the durations and fluences of the complete GRB sample follow Benford’s law. It is worth emphasizing that the higher the -value, the more likely the null hypothesis. In our present study, we exclude a null hypothesis if (equivalent to under the 95% confidence level).
GRBs can be divided into two classes, long GRBs and short GRBs, with a division line at s (Kouveliotou et al. 1993). Using the conventional division between the long and short GRB groups ( s and s, respectively), we find that there were long GRBs and short GRBs in the Fermi-GBM burst catalog. In this work, we also check if the durations and fluences of long and short GRBs obey Benford’s law. The first digit distributions of the duration and fluence of long and short GRBs are depicted in Figures 2 and 3, respectively. As listed in Table 1, the tests and the corresponding -values for the fluences are extremely supportive to the null hypothesis that the fluences of both long and short GRBs conform to Benford’s law. However, the durations of both long and short GRBs obviously deviate from Benford’s law. We find that the Pearson for the durations of long GRBs and short GRBs are and , which correspond to -values of 0.01 and , respectively. On the basis of the -values, we can safely reject the null hypothesis.
Similarly, in Figure 4 and Table 1, we show the first digit distribution of the energy flux of available GRBs. As we can see that the relative rank of the probability of occurrence of leading digits roughly agrees with Benford’s law, but the Benford distribution is not scrupulously obeyed. The rather large value and the extremely low value suggest that the energy flux of GRBs does not adhere to Benford’s law.
| Dataset | Physical quantity | Number | -value | |
| All GRBs | Duration | 3364 | 13.0/8 | 0.11 |
| Fluence | 3364 | 12.5/8 | 0.13 | |
| Long GRBs | Duration | 3061 | 19.9/8 | 0.01 |
| Fluence | 3061 | 11.5/8 | 0.18 | |
| Short GRBs | Duration | 603 | 74.2/8 | |
| Fluence | 603 | 7.1/8 | 0.53 | |
| All GRBs | Energy flux | 2298 | 67.4/8 |
3 Summary
In this work, we have performed a systematic investigation on the first digit distributions of the duration, fluence, and energy flux of GRBs. For the complete GRB sample, our results show that the first digits of the duration and fluence are not uniformly distributed, but that small digits are more common than big ones according to a logarithmic distribution as expected by Benford’s law. However, not all quantities obey Benford’s law. Artificial and restricted quantities often deviate from the law, such as the energy fluxes of GRBs in our study. The main reason for the deviation may be that the energy flux measurements are restricted by spectral fits for each GRB, and only 63% of bursts have the required spectral information. Thus, there is not enough dynamic range to make the energy fluxes be fully compliant with Benford’s law.
The Fermi-GBM burst catalog also revealed a bimodal distribution, and two classes of GRBs, i.e., long versus short GRBs with a separation at about 2 s, were confirmed (Kouveliotou et al. 1993; von Kienlin et al. 2020). Using the conventional long–short GRB classification scheme, we found that there were long GRBs (with s) and short GRBs (with s). Here, the first digits of the duration and fluence of long and short GRBs were also examined for adherence to Benford’s distribution. We found that the fluences of both long and short GRBs still follow Benford’s law, but their durations are no longer consistent with this law. Our results indicate that the data of fluence seems to be very natural and believable, but that is not always a good quantity to conduct GRB classification. That is, the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme.
Acknowledgements.
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Natural Science Foundation of China (grant Nos. 12373053 and 12321003), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).References
- Alexopoulos & Leontsinis (2014) Alexopoulos, T., & Leontsinis, S. 2014, Journal of Astrophysics and Astronomy, 35, 639
- Band et al. (1993) Band, D., Matteson, J., Ford, L., et al. 1993, ApJ, 413, 281
- Barlow & Bareiss (1985) Barlow, J. L., & Bareiss, E. H. 1985, Computing, 34, 325347
- Benford (1938) Benford, F. 1938, Proceedings of the American Philosophical Society, 78, 551
- Berger & Hill (2007) Berger, A., & Hill, T. P. 2007, The American Mathematical Monthly, 114, 588
- Berger et al. (2008) Berger, A., Hill, T. P., & Morrison, K. E. 2008, Journal of Theoretical Probability, 21, 91
- Buck et al. (1993) Buck, B., Merchant, A. C., & Perez, S. M. 1993, European Journal of Physics, 14, 59
- Burgos & Santos (2021) Burgos, A., & Santos, A. 2021, American Journal of Physics, 89, 851
- Burke & Kincanon (1991) Burke, J., & Kincanon, E. 1991, American Journal of Physics, 59, 952
- Cáceres et al. (2008) Cáceres, J. L. H., García, J. L. P., Martínez, C. M., Ortiz, & Domínguez, L. G. 2008
- Dantuluri & Desai (2018) Dantuluri, A., & Desai, S. 2018, Physica A Statistical Mechanics and its Applications, 506, 919
- De Ceuster et al. (1998) De Ceuster, M. J. K., Dhaene, G., & Schatteman, T. 1998, Journal of Empirical Finance, 5, 263
- de Jong et al. (2020) de Jong, J., de Bruijne, J., & De Ridder, J. 2020, A&A, 642, A205
- Eichler et al. (1989) Eichler, D., Livio, M., Piran, T., & Schramm, D. N. 1989, Nature, 340, 126
- Geyer & Williamson (2004) Geyer, C. L., & Williamson, P. P. 2004, Communications in Statistics - Simulation and Computation, 33, 229
- Gruber et al. (2014) Gruber, D., Goldstein, A., Weller von Ahlefeld, V., et al. 2014, ApJS, 211, 12
- Hill (1995) Hill, T. P. 1995, Statistical Science, 10, 354
- Kouveliotou et al. (1993) Kouveliotou, C., Meegan, C. A., Fishman, G. J., et al. 1993, ApJ, 413, L101
- Ley (1996) Ley, E. 1996, The American Statistician, 50, 311
- Mamidipaka & Desai (2023) Mamidipaka, P., & Desai, S. 2023, Astroparticle Physics, 144, 102761
- Melita & Miraglia (2021) Melita, M. D., & Miraglia, J. E. 2021, New A, 89, 101654
- Moret et al. (2006) Moret, M. A., de Senna, V., Pereira, M. G., & Zebende, G. F. 2006, International Journal of Modern Physics C, 17, 1597
- Narayan et al. (1992) Narayan, R., Paczynski, B., & Piran, T. 1992, ApJ, 395, L83
- Narayana Bhat et al. (2016) Narayana Bhat, P., Meegan, C. A., von Kienlin, A., et al. 2016, ApJS, 223, 28
- Newcomb (1881) Newcomb, S. 1881, American Journal of Mathematics, 4, 39
- Ni et al. (2009) Ni, D.-D., Wei, L., & Ren, Z.-Z. 2009, Communications in Theoretical Physics, 51, 713
- Ni & Ren (2008) Ni, D., & Ren, Z. 2008, European Physical Journal A, 38, 251
- Nigrini (1996) Nigrini, M. J. 1996, J. Am. Tax. Assoc., 18, 72
- Nigrini & Mittermaier (1997) Nigrini, M. J., & Mittermaier, L. J. 1997, Auditing-a Journal of Practice & Theory, 16, 52
- Paczyński (1998) Paczyński, B. 1998, ApJ, 494, L45
- Pinkham (1961) Pinkham, R. S. 1961, Annals of Mathematical Statistics, 32, 1223
- Schatte (1988) Schatte, P. 1988, J. Inf. Process. Cybern., 24, 443
- Shao & Ma (2009) Shao, L., & Ma, B.-Q. 2009, Modern Physics Letters A, 24, 3275
- Shao & Ma (2010a) Shao, L., & Ma, B.-Q. 2010a, Astroparticle Physics, 33, 255
- Shao & Ma (2010b) Shao, L., & Ma, B.-Q. 2010b, Physica A Statistical Mechanics and its Applications, 389, 3109
- Shukla et al. (2017) Shukla, A., Pandey, A. K., & Pathak, A. 2017, Journal of Astrophysics and Astronomy, 38, 7
- Sottili et al. (2012) Sottili, G., Palladino, D. M., Giaccio, B., & Messina, P. 2012, Mathematical Geosciences, 44, 619
- von Kienlin et al. (2014) von Kienlin, A., Meegan, C. A., Paciesas, W. S., et al. 2014, ApJS, 211, 13
- von Kienlin et al. (2020) von Kienlin, A., Meegan, C. A., Paciesas, W. S., et al. 2020, ApJ, 893, 46
- Woosley & Bloom (2006) Woosley, S. E., & Bloom, J. S. 2006, ARA&A, 44, 507